OpenStudy (anonymous):

which ordered pair is a solution of the equation y=3x (-2,-9) (-8,-18) (-8,-3) (-10,-30) which ordered pair is a solution of the equation y=-9x + 4 (10,-86) (-4, -58) (6,-41) (-6, 57) please help explain these to me

OpenStudy (anonymous):

(-10,-30) (10,-86)

OpenStudy (anonymous):

Substitute the values into the equations. Let's check @surdawi answers: 1) y=3x (-10, -30) -30 = 3(-10) -30 = -30 *CORRECT* 2) y=-9x + 4 (10,-86) -86 = -9(10) + 4 -86 = -90 + 4 - 86 = -86 *CORRECT*